OpenStudy (anonymous):

@jim_thompson5910 Could you please help me with this problem?

OpenStudy (anonymous):

I think it would be 6*4 = 24 ?

jimthompson5910 (jim_thompson5910):

each side of the shaded square is not 6

OpenStudy (anonymous):

It is 20

jimthompson5910 (jim_thompson5910):

to get the side length x, you need to solve for x 6/(9+6) = x/15 oh I get x = 6, so you were right before in saying 6*4 = 24 I had the wrong set up at first

OpenStudy (anonymous):

Yeah, I'm saying for my final answer I got 24.

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (anonymous):

My options were : 24 cm 36 cm 12 cm 42 cm and I got 24.

jimthompson5910 (jim_thompson5910):

it's correct

OpenStudy (anonymous):

Thanks!